OpenStudy (anonymous):

Solve equation. Convert improper fractions to mixed or proper fractions. 7 3/5 + 8 2/3

OpenStudy (anonymous):

20/15 or 1/3 ? right ?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

mulitply 5 and 7 then add 3 to get the first mixed fraction multiply 3 and 8 then add 2 to get the second, then add the two mixed fractions using a common denomintaor.

OpenStudy (anonymous):

tell me what you get when you are done :)

OpenStudy (anonymous):

i already typed my answer before you explained it to me so can you tell me if im right or not please ?

OpenStudy (anonymous):

@lgbasallote : is my answer right ?

OpenStudy (anonymous):

@bronzegoddess

OpenStudy (lgbasallote):

my calculator disagrees lol

OpenStudy (anonymous):

o okay, lemme work it ou then :)

OpenStudy (lgbasallote):

the goddess will explain (im sleeping now) <tips hat>

OpenStudy (anonymous):

44/15 ?

OpenStudy (lgbasallote):

nope..BIGGER...a lot

OpenStudy (lgbasallote):

7 3/5 = 38/5

OpenStudy (lgbasallote):

8 2/3 = 26/3

OpenStudy (lgbasallote):

okay seriously goddess will explain =)) im out now

OpenStudy (anonymous):

so 64/15 ? can that be simplified ?

OpenStudy (anonymous):

@bronzegoddess

OpenStudy (anonymous):

(3 x 38 ) + (5x26) =

OpenStudy (anonymous):

your numerator is wrong.

OpenStudy (anonymous):

it should be 244/15

OpenStudy (anonymous):

@Lil-Bit012 do you understand?

OpenStudy (anonymous):

@Lil-Bit012